40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 28 02:43:49 2013, yuta, Update, LSC, some qualitative evidence of PRMI sideband lock PRMIsideband.png
    Reply  Thu Feb 28 03:25:27 2013, yuta, Update, LSC, PR gain ~ 25 from PRMI carrier lock PRMIcarrier.png
       Reply  Thu Feb 28 03:41:31 2013, Koji, Update, LSC, PR gain ~ 25 from PRMI carrier lock 
       Reply  Sat Mar 2 05:53:15 2013, yuta, Update, LSC, stable lock of PRMI LSCMICHOLTF_PRMI.pngLSCPRCLOLTF_PRMI.pngPRMtoAS55REFL55.png
          Reply  Sat Mar 2 14:52:02 2013, rana, Update, LSC, stable lock of PRMI 
             Reply  Sun Mar 3 22:16:46 2013, Jenne, Update, LSC, LSC whitening triggering started 
                Reply  Mon Mar 4 09:55:33 2013, rana, Update, LSC, LSC whitening triggering started 
                Reply  Tue Mar 5 18:36:27 2013, Jenne, Update, LSC, LSC whitening triggering started 
Message ID: 8215     Entry time: Sun Mar 3 22:16:46 2013     In reply to: 8213     Reply to this: 8217   8234
Author: Jenne 
Type: Update 
Category: LSC 
Subject: LSC whitening triggering started 

[Jenne, Annalisa]

We have started working on writing the c-code to parse the LSC input matrix, to see which PD is used for what degree of freedom, and to output a trigger for the PD.  The code is in ..../isc/c1/src, and there is a little block in the LSC code to the left of the triggering stuff.  Right now, the output of the c-code just goes to some temporary EPICS channels, so that we can see if things are working before we actually implement it.  At this time, there is no change to how the LSC model runs.

I can't figure out a bug in my c-code though.  Right now it's all commented out, so that the LSC model would start, but if I try to sum all of the elements in an array, the model compiles fine, but it won't start running.  I'm going to ask Jamie about it tomorrow.  I have a less-tidy backup plan if we can't get this figured out.

If I have time on the IFO to check that this works tomorrow, I expect another few hours of work (2?  3?), and then we'll have whitening filter triggering.

ELOG V3.1.3-